Joe Kernen of CNBC recently commented that Amazon (AMZN) will start paying dividends in the medium-term
future because of all of the cash flow that it generates (coupled with a cash balance in excess of $43 billion).
When pressed on the issue, Joe made a bold forecast that Amazon will start paying its first dividend 5 years from now
in the amount of $93.00 per share. Joe also projected that the dividends would grow, over the following 5 years, by
14.00% per year, after which the growth rate would be a constant rate of 7.00%/year, forever. If the appropriate
required return (also called discount rate, because we use it for discounting!) for AMZN's stock is 11.0%, what should
today’s stock price be if your calculation is to be based on Kernen’s projected information?
Answer #1

Dividend year 5 (D5) =93
D6 =D5*(1+g)^1=93*1.14
D7 =93*1.14^2
D8 =93*1.14^3
D9 =93*1.14^4
D10 =93*1.14^5
Terminal Value =93*1.14^5*(1+7%)/(11%-7%) =4789.9501

Price of Stock =93/(1+11%)^5+93*1.14/(1+11%)^6+93*1.14^2/(1+11%)^7+93*1.14^3/(1+11%)^8+93*1.14^4/(1+11%)^9+93*1.14^5/(1+11%)^10+4789.9501/1.11^10 =2041.29